what is the answer to -3+1/2-2 1/2
Answer by ptaylor(2198)

what is the answer to -3+1/2-2 1/2
-3+1/2-2 1/2=
-9/2+1/2-5/2=
-14/2+1/2=
-13/2=
-6 1/2-------------------------------------------answer
Hope this helps-------ptaylor